OpenStudy (anonymous):

Check my work please? If I have 7 men and 9 women, and I am making a guest list that will consist of 5 men and 5 women. How many different ways can I chose who I invite? I think that for the men, its 7 choose 5, and for the women 9 choose 5. Since order doesn't matter, I think that I just add these two right?

OpenStudy (anonymous):

so I have \[\left(\begin{matrix}7 \\ 5\end{matrix}\right) + \left(\begin{matrix}9 \\ 5 \end{matrix}\right)\] which is 147.

OpenStudy (anonymous):

Actually, you multiply those

OpenStudy (anonymous):

ah shoot, basic principle of counting eh?

OpenStudy (anonymous):

So, 2646 unique lists.

OpenStudy (anonymous):

The reasoning is that for every grouping of one gender, you have available the full gamut of all the other gender's possibilities.

OpenStudy (anonymous):

Yes, you got it! Good job!

OpenStudy (anonymous):

All good now, @xartaan ?

OpenStudy (anonymous):

Ya thanks, very clear, I appreaciate it!

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @xartaan